The sum of the digits of a two-digit number is 9. If the digits are reversed, the new number is 27 less than the original. Find the original number.
Answer by fanism(39)

assume the first digit is x and second digit is y
Since the sum of the digits of a two-digit number is 9, therefore
x + y = 9 -- eq1
it says if the digits are reversed, the new number is 27 less than the original.
since we are looking at the number like xy, to seperate them, it is actually 10x+y for x is a tens digit.
10y + x = 10x + y + 27
simplify it, u will get:
9y = 9x + 27
y = x + 3 -- eq2
sub eq2 into eq1, u will have
x + (x + 3) = 9
2x + 3 = 9
2x = 6
x = 3
put back into eq1,
3 + y = 9
y = 6
The original number is 36.